OpenStudy (anonymous):

f(x)=2x^2-7 why the Rolle's theorem don't apply

OpenStudy (amistre64):

first, define rolles thrm ....

OpenStudy (amistre64):

then check the problem and make sure you havent posted a mistake ... if that fails, take a screenshot and attach a picture of it to make sure theres no wonky stuff going on in translation ....

OpenStudy (anonymous):

I equal f(x) to zero and I have square root of 7/2. f prime gives me zero

OpenStudy (amistre64):

not quite sure how that relates ... rolles thrm applies to your stated problem unless youve mistyped something ... and that response is not readable to me

OpenStudy (anonymous):

I forgot something it has an interval between [0,4]

OpenStudy (amistre64):

so, we want to find 2 points a,b in the interval such that the slope from a to b is 0 \[\frac{f(a)-f(b)}{a-b}=0\] \[\frac{2a^2-7-(2b^2-7)}{a-b}=0\] \[\frac{2a^2-7-2b^2+7}{a-b}=0\] \[\frac{2a^2-2b^2}{a-b}=0\] \[\frac{2(a^2-b^2)}{a-b}=0\] so, this is only possible when a^2 - b^2 = 0 giving us a different of squares to factor into (a+b) or (a-b) = 0; so a=-b, or a=b fit the bill since the denominator is a-b, then a=b is out

OpenStudy (amistre64):

is there 2 number in the interval 0 to 4 such that a+b=0?

OpenStudy (anonymous):

thank you very much for your help but i don't understand this!

OpenStudy (amistre64):

you are looking for 2 number in the interval that have a slope of zero between them .... there are none in that the stated interval therefore Rolles thrm fails. Rolles thrm: if 2 points in an interval create a slope of 0 ... f(a)=f(b), then there must be a f'(c) = 0 for some c inbetween a and b

OpenStudy (amistre64):

since the interval is from 0 to 4, there are no additive inverses that can be used.

OpenStudy (anonymous):

so from the original equation you must have an x= to two numbers?

OpenStudy (amistre64):

2 different numbers yes

OpenStudy (anonymous):

thanks for being so patience with me!

OpenStudy (amistre64):

yep .. you never use Rolles thrm but its just there to prove the underling math